7X over 9 plus 3 equals one-eighth.
Answer by stanbon(75887)

7X over 9 plus 3 equals one-eighth.
-----------
(7x/9) + 3 = (1/8)
(7x/9) = -23/8
Cross multiply to get:
56x = -207
x = -207/56
==============
Cheers,
Stan H.